# Improving Autoformalization Using Direct Dependency Retrieval

### v3

@@ 3.1 Direct Dependency Retrieval
Difficulty Num Depend Rate Depend Length 0 835 0.202 395 209 580 838 310.202\,395\,209\,580\,838\,31 0.287 425 149 700 598 80.287\,425\,149\,700\,598\,8 1 120040 0.560 421 526 157 947 40.560\,421\,526\,157\,947\,4 1.074 716 761 079 640 21.074\,716\,761\,079\,640\,2 2 113816 0.785 056 582 554 298 10.785\,056\,582\,554\,298\,1 1.870 738 736 205 805 91.870\,738\,736\,205\,805\,9 3 49841 0.808 330 490 961 256 80.808\,330\,490\,961\,256\,8 2.088 682 008 787 9462.088\,682\,008\,787\,946 4 69882 0.795 927 420 508 857 80.795\,927\,420\,508\,857\,8 1.963 123 551 129 046 11.963\,123\,551\,129\,046\,1 5 67547 0.524 064 725 302 3820.524\,064\,725\,302\,382 1.112 691 903 415 399 71.112\,691\,903\,415\,399\,7 6 41312 0.813 032 532 920 216 90.813\,032\,532\,920\,216\,9 2.251 839 659 178 9312.251\,839\,659\,178\,931 7 36675 0.756 264 485 344 240.756\,264\,485\,344\,24 2.135 460 122 699 386 32.135\,460\,122\,699\,386\,3 8 8682 0.821 354 526 606 772 70.821\,354\,526\,606\,772\,7 2.430 085 233 817 0932.430\,085\,233\,817\,093 9 708 0.830 508 474 576 271 20.830\,508\,474\,576\,271\,2 2.632 768 361 581 9212.632\,768\,361\,581\,921
@@ 4 Direct Dependency Retrieval-augmented Autoformalization
Figure 3: LLM consistency across multiple attempts. Subplots (a–e) correspond to the performance of DeepSeek-R1 on datasets Diff01, Diff23, Diff45, Diff67, and Diff89, respectively. Each subplot illustrates the proportion of problems ($y$ axis) for which at least k (from 1 to 8, $x$ axis) out of 8 attempts successfully pass the BEq verification. Blue: DDR; Orange: ICL; Green: N/A.
@@ 4.1 Dataset Overview and Difficulty Analysis
The original difficulty ratings for the FineLeanCorpus (peng2025criticlean) datasets, assigned by an LLM grader, exhibit a notable discrepancy with our experimental results in Table 4 and Table 5. To investigate this inconsistency, we propose assessing the actual formalization difficulty using two alternative metrics: Depend Rate and Depend Length. Based on the data in Table 1, we can calculate the weighted averages of Depend Rate and Depend Length for the Diff23, Diff45, Diff67, and Diff89 subsets. The results are as follows: Diff45 has the lowest values (0.662, 1.545), the values for Diff23 (0.792, 1.937) and Diff67 (0.786, 2.197) are higher, and Diff89 has the highest values (0.822, 2.445). These metrics reveal a difficulty ranking of Diff89 > Diff67 > Diff23 > Diff45. This ranking aligns closely with our experimental observations, suggesting that Depend Rate and Depend Length, derived via our DDR and SAC, can serve as a potential indicator of formalization difficulty along with LLM-based graders. The Diff01 subset is excluded from this analysis because its imprecise informal descriptions introduce extraneous formalization challenges, making its difficulty hard to assess with these metrics. Detailed discussion are presented in appendix B
Figure 4: LLM efficiency across multiple attempts. Subplots (a–e) correspond to the performance of DeepSeek-R1 on datasets Diff01, Diff23, Diff45, Diff67, and Diff89, respectively. Each subplot illustrates the intermediate pass@k score ($y$ axis) out of 8 attempts using BEq verification, k range from 1 to 8 ($x$ axis). Blue: DDR; Orange: ICL; Green: N/A.
Method    Diff01    Diff23    Diff45    Diff67    Diff89 TypeC BEq TypeC BEq TypeC BEq TypeC BEq TypeC BEq MMA $0.47$ $0.03$ $0.455$ $0.075$ $0.6$ $0.125$ $0.5$ $0.045$ $0.385$ $0.03$ RA $0.67$ $0.04$ $0.74$ $0.13$ $0.77$ $0.22$ $0.73$ $0.085$ $0.61$ $0.03$ RA+R@5 0.880 0.070 0.900 $0.125$ $0.895$ $0.21$ $0.855$ $0.095$ $0.765$ $0.05$ RA+R@10 $0.815$ $0.045$ $0.86$ $0.115$ $0.84$ $0.215$ $0.825$ $0.085$ $0.73$ $0.055$ RA+DDR $0.87$ 0.070 0.900 0.140 0.930 0.255 0.865 0.105 0.810 0.065
@@ 4.3 Hallucination Study
Method RAG    Diff01    Diff23    Diff45    Diff67    Diff89 TypeC BEq TypeC BEq TypeC BEq TypeC BEq TypeC BEq Claude-3.5-Sonnet N/A $0.845$ $0.14$ $0.755$ $0.21$ 0.800 $0.275$ 0.680 0.130 $0.53$ $0.035$ Claude-3.5-Sonnet ICL $0.74$ $0.125$ $0.515$ $0.15$ $0.65$ $0.24$ $0.46$ $0.075$ $0.385$ $0.04$ Claude-3.5-Sonnet DDR 0.905 0.165 0.790 0.235 $0.77$ 0.305 $0.63$ 0.130 0.635 0.055 DeepSeek-R1 N/A $0.91$ $0.1$ $0.745$ $0.125$ $0.82$ $0.27$ $0.78$ $0.115$ $0.66$ $0.085$ DeepSeek-R1 ICL $0.885$ $0.11$ $0.76$ $0.22$ $0.845$ $0.34$ $0.76$ $0.145$ $0.575$ $0.08$ DeepSeek-R1 DDR 0.940 0.170 0.875 0.280 0.950 0.395 0.875 0.180 0.705 0.090 GPT-4o N/A $0.905$ $0.1$ $0.765$ 0.180 $0.8$ $0.175$ $0.75$ $0.09$ $0.575$ $0.05$ GPT-4o ICL $0.88$ $0.1$ $0.675$ $0.14$ $0.73$ $0.185$ $0.63$ $0.06$ $0.405$ $0.05$ GPT-4o DDR 0.940 0.125 0.775 0.180 0.850 0.280 0.780 0.100 0.630 0.060 Ling-flash-2.0 N/A 0.855 $0.105$ 0.595 0.150 0.670 $0.235$ $0.51$ $0.09$ $0.38$ 0.055 Ling-flash-2.0 ICL $0.775$ $0.075$ $0.42$ $0.1$ $0.6$ $0.225$ $0.395$ $0.08$ $0.295$ 0.055 Ling-flash-2.0 DDR $0.775$ 0.115 0.595 0.150 $0.64$ 0.260 0.535 0.105 0.440 0.055 Qwen-Max N/A $0.795$ $0.095$ $0.605$ 0.155 0.675 0.270 $0.62$ $0.11$ $0.44$ $0.060$ Qwen-Max ICL $0.73$ $0.065$ $0.445$ $0.105$ $0.555$ $0.17$ $0.445$ $0.075$ $0.35$ $0.060$ Qwen-Max DDR 0.835 0.110 0.665 $0.15$ $0.575$ $0.21$ 0.695 0.115 0.545 0.065
@@ 4.5 Autoformalization Results
To evaluate DDR’s potential as a plug-and-play component, we integrated it with ICL methods. Table 5 compares three experimental setups: (1) direct autoformalization, (2) two-stage autoformalization using a prompt-based dependency retrieval, and (3) two-stage autoformalization using a DDR-based dependency retrieval. The results indicate that the DDR-based RAG is the superior strategy in most test scenarios. Experiments on Deepseek-R1, illustrated in Figure 3, further reveal DDR’s stability advantage: the DDR-integrated method not only achieves a higher total pass@8 success count but also exhibits a significantly better distribution of successful attempts across 8 independent trials compared to the baseline. Furthermore, Figure 4 shows that this method can complete more tasks with fewer attempts.
@@ A.1 Full Figures
Figure 5: LLM consistency across multiple attempts. From left to right correspond to Diff01, Diff23, Diff45, Diff67, and Diff89; from top to bottom correspond to Claude-3.5-Sonnet, DeepSeek-R1, GPT-4o, Ling-flash-2.0, and Qwen-Max. Each subplot illustrates the proportion of problems ($y$ axis) for which at least k (from 1 to 8, $x$ axis) out of 8 attempts successfully pass the BEq verification. Blue: DDR; Orange: ICL; Green: N/A.
Figure 6: LLM efficiency across multiple attempts. From left to right correspond to Diff01, Diff23, Diff45, Diff67, and Diff89; from top to bottom correspond to Claude-3.5-Sonnet, DeepSeek-R1, GPT-4o, Ling-flash-2.0, and Qwen-Max. Each subplot illustrates the intermediate pass@k score ($y$ axis) out of 8 attempts using BEq verification, k range from 1 to 8 ($x$ axis). Blue: DDR; Orange: ICL; Green: N/A.
@@ C.4 Baseline Details
For a fair comparison, we adopt an experimental setup identical to that of liu2025rethinking. We used the same prompt as described in Section A.7.1 of liu2025rethinking for evaluation on BEq. Similarly, the prompt for the ICL method is also consistent with that in Section A.7.3 of liu2025rethinking. DeepSeek-Prover-v2 (ren2025deepseek) is utilized as an auxiliary prover for BEq.
@@ D.3 Other Templates
Other ICL templates employed in this experiment, including those for autoformalization without dependency retrieval and BEq verification, adhere to the corresponding settings in liu2025rethinking.

### v4

@@ 3.1 Direct Dependency Retrieval
Difficulty Num Depend Rate Depend Length 0 835 0.188 023 952 095 808 380.188\,023\,952\,095\,808\,38 0.267 065 868 263 473 060.267\,065\,868\,263\,473\,06 1 120040 0.555 081 639 453 515 50.555\,081\,639\,453\,515\,5 1.061 096 301 232 922 41.061\,096\,301\,232\,922\,4 2 113816 0.783 756 238 138 750 30.783\,756\,238\,138\,750\,3 1.862 936 669 712 518 51.862\,936\,669\,712\,518\,5 3 49841 0.807 347 364 619 490.807\,347\,364\,619\,49 2.079 312 212 836 8212.079\,312\,212\,836\,821 4 69882 0.795 240 548 352 937 80.795\,240\,548\,352\,937\,8 1.957 700 123 064 594 51.957\,700\,123\,064\,594\,5 5 67547 0.523 383 718 003 760 40.523\,383\,718\,003\,760\,4 1.105 689 371 844 789 51.105\,689\,371\,844\,789\,5 6 41312 0.810 878 195 197 521 40.810\,878\,195\,197\,521\,4 2.229 908 985 282 726 52.229\,908\,985\,282\,726\,5 7 36675 0.753 946 830 265 848 60.753\,946\,830\,265\,848\,6 2.108 166 325 835 037 42.108\,166\,325\,835\,037\,4 8 8682 0.819 511 633 264 224 80.819\,511\,633\,264\,224\,8 2.393 112 186 132 227 82.393\,112\,186\,132\,227\,8 9 708 0.824 858 757 062 146 80.824\,858\,757\,062\,146\,8 2.611 581 920 903 9552.611\,581\,920\,903\,955
@@ 4 Direct Dependency Retrieval-augmented Autoformalization
Figure 3: LLM consistency across multiple attempts. Subplots (a–e) correspond to the performance of DeepSeek-R1 on datasets Diff01, Diff23, Diff45, Diff67, and Diff89, respectively. Each subplot illustrates the proportion of problems ($y$ axis) for which at least k (from 1 to 8, $x$ axis) out of 8 attempts successfully pass the BEq verification. Dependency retrieval method: DDR (blue); ICL (orange); N/A (green).
@@ 4.1 Dataset Overview and Difficulty Analysis
The original difficulty ratings for the FineLeanCorpus (peng2025criticlean) datasets, assigned by an LLM grader, exhibit a notable discrepancy with our experimental results in Table 4 and Table 5. To investigate this inconsistency, we propose assessing the actual formalization difficulty using two alternative metrics: Depend Rate and Depend Length. Based on the data in Table 1, we can calculate the weighted averages of Depend Rate and Depend Length for the Diff23, Diff45, Diff67, and Diff89 subsets. The results are as follows: Diff45 has the lowest values (0.662, 1.539), the values for Diff23 (0.791, 1.929) and Diff67 (0.784, 2.173) are higher, and Diff89 has the highest values (0.820, 2.410). These metrics reveal a difficulty ranking of Diff89 > Diff67 > Diff23 > Diff45. This ranking aligns closely with our experimental observations, suggesting that Depend Rate and Depend Length, derived via our DDR and SAC, can serve as a potential indicator of formalization difficulty along with LLM-based graders. The Diff01 subset is excluded from this analysis because its imprecise informal descriptions introduce extraneous formalization challenges, making its difficulty hard to assess with these metrics. Detailed discussion are presented in appendix B
Figure 4: LLM efficiency across multiple attempts. Subplots (a–e) correspond to the performance of DeepSeek-R1 on datasets Diff01, Diff23, Diff45, Diff67, and Diff89, respectively. Each subplot illustrates the intermediate pass@k score ($y$ axis) out of 8 attempts using BEq verification, k range from 1 to 8 ($x$ axis). Dependency retrieval method: DDR (blue); ICL (orange); N/A (green).
Method    Diff01    Diff23    Diff45    Diff67    Diff89 TypeC BEq TypeC BEq TypeC BEq TypeC BEq TypeC BEq MMA $0.47$ $0.095$ $0.455$ $0.11$ $0.6$ $0.19$ $0.5$ $0.1$ $0.385$ $0.065$ RA $0.67$ $0.115$ $0.730$ $0.170$ $0.765$ $0.295$ $0.735$ $0.185$ $0.61$ $0.105$ RA+R@5 $0.86$ $0.185$ $0.795$ $0.18$ $0.805$ $0.145$ $0.825$ $0.115$ $0.765$ $0.135$ RA+R@10 0.880 $0.205$ $0.86$ $0.2$ 0.910 $0.255$ $0.825$ $0.16$ $0.765$ 0.150 RA+DDR $0.865$ 0.235 0.900 0.215 0.910 0.305 0.865 0.225 0.770 0.150
@@ 4.3 Hallucination Study
Method RAG    Diff01    Diff23    Diff45    Diff67    Diff89 TypeC BEq TypeC BEq TypeC BEq TypeC BEq TypeC BEq Claude-3.5-Sonnet N/A $0.83$ $0.415$ $0.74$ $0.33$ 0.775 $0.385$ 0.680 $0.205$ $0.53$ $0.1$ Claude-3.5-Sonnet ICL $0.74$ $0.37$ $0.515$ $0.245$ $0.64$ $0.33$ $0.455$ $0.145$ $0.385$ $0.11$ Claude-3.5-Sonnet DDR 0.905 0.460 0.770 0.335 $0.765$ 0.425 $0.625$ 0.220 0.630 0.130 DeepSeek-R1 N/A $0.91$ $0.225$ $0.745$ $0.205$ $0.82$ $0.36$ $0.78$ $0.23$ $0.66$ $0.13$ DeepSeek-R1 ICL $0.88$ $0.34$ $0.755$ $0.315$ $0.825$ $0.44$ $0.76$ $0.27$ $0.57$ $0.135$ DeepSeek-R1 DDR 0.940 0.410 0.875 0.385 0.940 0.515 0.875 0.345 0.695 0.200 GPT-4o N/A $0.905$ $0.44$ $0.76$ $0.295$ $0.8$ $0.4$ $0.745$ $0.24$ $0.575$ $0.135$ GPT-4o ICL $0.88$ 0.460 $0.675$ $0.255$ $0.73$ $0.345$ $0.625$ $0.21$ $0.405$ $0.12$ GPT-4o DDR 0.940 0.460 0.775 0.320 0.850 0.455 0.780 0.275 0.630 0.150 Ling-flash-2.0 N/A 0.850 0.405 0.610 $0.21$ 0.665 $0.3$ $0.51$ $0.19$ $0.38$ $0.085$ Ling-flash-2.0 ICL $0.775$ $0.355$ $0.42$ $0.17$ $0.59$ $0.28$ $0.395$ $0.17$ $0.295$ $0.085$ Ling-flash-2.0 DDR $0.775$ $0.4$ $0.595$ 0.225 $0.63$ 0.360 0.535 0.205 0.440 0.115 Qwen-Max N/A $0.795$ 0.375 $0.605$ $0.22$ 0.665 0.360 $0.62$ $0.205$ $0.44$ $0.105$ Qwen-Max ICL $0.73$ $0.315$ $0.445$ $0.16$ $0.55$ $0.26$ $0.445$ $0.165$ $0.35$ $0.115$ Qwen-Max DDR 0.830 0.375 0.660 0.250 $0.56$ $0.285$ 0.695 0.255 0.545 0.120
@@ 4.5 Autoformalization Results
To evaluate DDR’s potential as a plug-and-play component, we integrated it with ICL methods. Table 5 compares three experimental setups: (1) direct autoformalization, (2) two-stage autoformalization using a prompt-based dependency retrieval, and (3) two-stage autoformalization using a DDR-based dependency retrieval. The results indicate that the DDR-based RAG is the superior strategy in most test scenarios. Experiments on Deepseek-R1, illustrated in Figure 3, further reveal DDR’s stability advantage: the DDR-integrated method not only achieves a higher total pass@8 success count but also exhibits a significantly better distribution of successful attempts across 8 independent trials compared to the baseline. Furthermore, Figure 4 shows that DDR can complete more tasks with fewer attempts.
@@ A.1 Full Figures
Figure 5: LLM consistency across multiple attempts. From left to right correspond to Diff01, Diff23, Diff45, Diff67, and Diff89; from top to bottom correspond to Claude-3.5-Sonnet, DeepSeek-R1, GPT-4o, Ling-flash-2.0, and Qwen-Max. Each subplot illustrates the proportion of problems ($y$ axis) for which at least k (from 1 to 8, $x$ axis) out of 8 attempts successfully pass the BEq verification. Dependency retrieval method: DDR (blue); ICL (orange); N/A (green).
Figure 6: LLM efficiency across multiple attempts. From left to right correspond to Diff01, Diff23, Diff45, Diff67, and Diff89; from top to bottom correspond to Claude-3.5-Sonnet, DeepSeek-R1, GPT-4o, Ling-flash-2.0, and Qwen-Max. Each subplot illustrates the intermediate pass@k score ($y$ axis) out of 8 attempts using BEq verification, k range from 1 to 8 ($x$ axis). Dependency retrieval method: DDR (blue); ICL (orange); N/A (green).
@@
A.2 Out of Domain Study
To evaluate the generalization ability of DDR, we conduct autoformalization experiments on the Omni-math dataset (gao2024omni), adopting the experimental setup from lin2025goedel. We compare DDR against the baseline methods Goedel-Formalizer-V2-8B (lin2025goedel) and Kimina-Autoformalizer-7B (kimina_prover_2025). To evaluate the correctness of the generated formalizations, we employ the method proposed in lin2025goedel, which combines Lean4 TypeC with semantic judgments from LLMs. The pass@8 results are presented in Table 6. The results demonstrate that DDR enhances the performance of existing autoformalizers, even those not specifically designed for dependency retrieval.
Name RAG Pass Failed Goedel-Formalizer-V2-8B N/A 180 120 Goedel-Formalizer-V2-8B DDR 183 117 Kimina-Autoformalizer-7B N/A 105 195 Kimina-Autoformalizer-7B DDR 128 172
Table 6: Comparison of different formalizers on 300 Omni-math problems.
@@ C.4 Baseline Details
For a fair comparison, we adopt an experimental setup identical to that of liu2025rethinking. We used the same prompt as described in Section A.7.1 of liu2025rethinking for evaluation on BEq. Similarly, the prompt for the ICL method is also consistent with that in Section A.7.3 of liu2025rethinking. Gemini 2.5 pro is utilized as an auxiliary tool for BEq. Gemini 2.5 pro is also the LLM judge used in section A.2.
@@ D.3 Other Templates
Other ICL templates employed in this paper, including those for autoformalization without dependency retrieval and BEq verification, adhere to the corresponding settings in liu2025rethinking.